OpenStudy (anonymous):

an electrician cuts a 136-ft piece of cable. one piece is 16 feet less than 3 times the length of the other piece. Find the length of each piece.

OpenStudy (anonymous):

x + (3x - 16) = 136 4x - 16 = 136 4x = 152 x = 38 feet the other piece is 3x - 16 = 3(38) - 16 = 114 - 16 = 98 feet.

OpenStudy (anonymous):

4x-16=136 ==> x=38 ==> the 2nd piece 136-38 =9m

OpenStudy (anonymous):

the 2nd piece is 98